# Revisiting Transformers with Insights from Image Filtering

### v1

@@ V-C Vision Tasks on ImageNet-1k and ADE20k
Table IV: Image Segmentation on ADE20k
@@
Empirical analysis. Additionally, in Figure 4, we visualize attention heatmaps of the last layer, derived from the normalized attention weights between the class token and the image patches averaged over heads, for both DeiT with $K_{SA}$ and DeiT Bilateral with $K_{BSA}$. We note that the latter attains more coverage of semantically important objects within the image sample. As we have tested on numerous samples (see Appendix VII), we hypothesize that this behavior, when happens, is partially influenced by the presence of two additional “rough” cross-terms (token-to-position and position-to-token) in the standard self-attention mechanism. As a result, most heads tend to focus on the region with the strongest positive signal, which can dominate these cross-term interactions, while $K_{BSA}$ is free of this distraction. Figure 5, on the other hand, shows the layer-wise average cosine similarities of the Transformer block output tokens, supporting the claim of Proposition IV.1.
Figure 4: Attention Heatmaps. Left to Right: Input, DeiT, and Deit-Bilateral.
VI Discussion on Framework Expressivity and Related Work
VI-A Attention Mechanisms and PE
The image filtering perspective provided in this work effectively builds upon and corrects similar notions initially taken by [12] and then [20] that overlook the presence of positional information. Consequently, the framework becomes more expressive and interpretable as we shall discuss next.
Notably, the empirical observations reported in [30] align closely with our theoretical perspective. The authors of [30] identify uniform and uninformative patterns in token-to-position proximity logits, suggesting the need for independent transformations of positional vectors to mitigate their influence. Essentially, their empirical observations resembles the Bilateral Attention formulation as the following form of logits are employed:
$\frac{\bm{x}^{\top}_{i}\bm{W}_{Q}^{\top}\bm{W}_{K}\bm{x}_{j}}{\sqrt{2d}}+\frac{\bm{p}^{\top}_{i}\bm{H}_{Q}^{\top}\bm{H}_{K}\bm{p}_{j}}{\sqrt{2d}}+b_{ij},$ (30)
where $\bm{H}_{Q}$ and $\bm{H}_{K}$ are independent learnable projections exclusively for PE, resembling the usage of different bandwidth for BF (Section III-B). Furthermore, Transformers with no positional encodings have also been observed to produce performance on par with absolute, relative and learnable positional encoding methods with somewhat better length generalization properties as studied by [31]. Our framework provides a theoretical justification for these empirical observations through the lens of NLM filtering (Section III-B), where we connect no positional encoding to a bilateral attention with infinitesimal weights given to positional weights.
Beyond offering a unifying perspective on self-attention, our work underscores the potential for leveraging insights from classical image processing techniques to inspire novel attention mechanisms. For instance, Elliptical Attention, a recently proposed variant in [25], computes attention scores as
$\alpha_{ij}=\mathrm{softmax}\left(\frac{\bm{q}_{i}^{\top}\bm{M}\bm{k}_{j}}{\sqrt{d}}\right),$ (31)
where $\bm{M}$ is a diagonal matrix with $M_{ii}$ representing the average function variability along the feature dimension $i$. This can be interpreted through the lens of anisotropic nonlocal means denoising techniques [45] and adaptive filtering methods [23], which construct spatially varying, non-spherical receptive fields to enhance feature preservation. Similarly, Twicing Attention [20] incorporates a residual correction technique from nonparametric regression, reminiscent of the twicing method in image processing, which our generalized framework naturally encompasses when additional computation is permitted as depicted in the following chart:
@@ VI-A Attention Mechanisms and PE
VI-B What about MLP blocks?
MLPs are usually not straightforward to interpret. The interpretability of MLPs often requires them to be wide-and-sparse [46]. Therefore, we shall discuss the sparse mixture of experts (SMoE) layer employed in most of the current state-of-the-art models [47, 48], conjecturing that its computation may bear a resemblance to techniques of sparse coding that are often utilized in k-SVD denoising [28].
In a SMoE layer, $M>1$ MLPs $f_{j}(\bm{x})=\bm{P}_{j}\sigma(\bm{Q}_{j}\bm{x})$, known as experts, are used instead of a single dense one to enhance the model capacity. Only a few of them are then activated for each input token conditionally. To be more precise, given an input token $\bm{x}\in\mathbb{R}^{d}$, the output of the SMoE layer is given by
$\bm{y}=\sum_{j=1}^{M}g_{j}(\bm{x},\bm{\theta})f_{j}(\bm{x})=\sum_{j=1}^{M}g_{j}(\bm{x},\bm{\theta})\bm{P}_{j}\sigma(\bm{Q}_{j}\bm{x}),$ (32)
@@ VI-B What about MLP blocks?
$g_{j}(\bm{x},\bm{\theta})=\begin{cases}\mathrm{softmax}(\bm{x}^{\top}\bm{\theta}_{j}),&\text{if }j\in\text{Top-}k(\bm{x}^{\top}\bm{\theta}_{i})\\ 0,&\text{otherwise}.\end{cases}$ (33)
Intuitively, the router functions as a means to cluster the input tokens by assigning groups of tokens to different experts by a similarity measure. This aspect of SMoE architecture has indeed been explored by a recent line of research [49, 50]. We also recall that clustering is a key technique behind most image segmentation algorithms as well [51, 52, 53].
Now to explore the conjecture mentioned in the beginning of this section, we can rewrite the output in a large matrix form, which highlights the sparse and structured nature of the computation, potentially mirroring the sparse coding principles [54] used in k-SVD denoising [28]. Following [55], let us define the intermediate output of each expert $j$ as $\bm{z}_{j}=\sigma(\bm{Q}_{j}\bm{x})\in\mathbb{R}^{k^{\prime}}$, where $k^{\prime}$ is the dimension of the intermediate representation. The contribution of expert $j$ to the output $\bm{y}$ is then $g_{j}(\bm{x},\bm{\theta})\bm{P}_{j}\bm{z}_{j}$, and the total output is the sum of these contributions over all $M$ experts.
Now, define $\bm{D}$ as the horizontal concatenation of the output transformation matrices $\bm{P}_{j}$ for all $M$ experts $\bm{D}=[\bm{P}_{1},\bm{P}_{2},\dots,\bm{P}_{M}]\in\mathbb{R}^{d\times(Mk)}$, where each $\bm{P}_{j}\in\mathbb{R}^{d\times k^{\prime}}$, so $\bm{D}$ has $d$ rows and $Mk^{\prime}$ columns. Next, construct a vector $\bm{z}\in\mathbb{R}^{Mk^{\prime}}$ by vertically stacking the scaled intermediate outputs $g_{j}(\bm{x},\bm{\theta})\bm{z}_{j}$ as $\bm{z}=\begin{bmatrix}g_{1}(\bm{x},\bm{\theta})\bm{z}_{1}&g_{2}(\bm{x},\bm{\theta})\bm{z}_{2}&\dots&g_{M}(\bm{x},\bm{\theta})\bm{z}_{M}\end{bmatrix}^{\top}$.
@@
With these definitions, the SMoE output can be expressed as a single matrix-vector product $\bm{y}=\bm{D}\bm{z}$ by construction. The matrix $\bm{D}$ can be viewed as a large “dictionary” of transformations, with $Mk$ columns, while the sparsity of $\bm{z}$ ensures that only a small subset of these transformations (corresponding to the $k$ active experts) is used for each input token. We conjecture that this formulation may be related to k-SVD denoising [54, 28], where a noisy signal is approximated using a sparse linear combination of dictionary atoms to produce a denoised output.
Table V: Forward Pass Efficiency in terms of FLOPs (GMac)
@@
VII Concluding Remarks
Summary. In this work, we developed a unifying image processing framework to better understand the feature-learning mechanism behind architectural components of Transformer-based models. In addition to addressing the frequently asked question of why positional encodings (PE) are added rather than concatenated in an investigative approach, we also explore how properly incorporating positional information enhances the accuracy and long-range capacity of the self-attention mechanism. Furthermore, we highlight the crucial role of residual connections in improving the model’s fidelity to its input and, as a result, robustness against adversarial attacks.
Future directions. Despite the growing interest in drawing inspiration from classical image processing for designing Transformer architectures [12, 20], several powerful techniques remain underexplored. In particular, while locally adaptive regression kernels (LARK) have been widely used in image denoising and feature extraction [23], more general variants [56] are yet to be systematically transferred to Transformers. Similarly, fast approximations of the bilateral filter [57], which enable efficient edge-preserving smoothing in real-time applications, could have been considered as a foundation for designing computationally efficient self-attention mechanisms.
Limitations. A specific limitation of Bilateral Attention proposed in this work is that it introduces an additional training computation complexity (note it does not pose any additional inference cost since positional embeddings are input-independent and can be precomputed after training, see Table V), which is undesirable under specific resource restricted scenarios. Moreover, a room of ambiguity still remains when it comes to understanding MLP blocks (if possible) other than the attempts in the previous section, which is an interesting open question for further exploration.
[Proofs and Extra Material]
More Attention Heatmaps
In Figure 6, we provide few more attention heatmaps averaged over attention heads for DeiT [6] and DeiT Bilateral. One can observe that the heatmaps produced by DeiT Bilateral usually captures extended meaningful regions of input samples which also implies a slightly better head diversity.
Figure 6: Left to Right: Input, DeiT, DeiT Bilateral.
Proof of Theorem III.1
Proof. The proof is similar to the derivation of Eqn. 12 but with vector valued intensity values. The signal $u(\bm{p})\in\mathbb{R}^{d}$ at position $\bm{p}$ is estimated using a (nonparametric) point estimation framework. Specifically, this involves solving a weighted least squares problem:
$\hat{u}(\bm{p}_{i})=\arg\min_{u(\bm{p}_{i})}\mathcal{J}_{K}(u(\bm{p}_{i})),$ (34)
@@ VII Concluding Remarks
$\hat{u}(\bm{p}_{i})=\frac{\sum_{j=1}^{n}\bm{y}_{j}K(\bm{p}_{i},\bm{p}_{j},\bm{y}_{i},\bm{y}_{j})}{\sum_{j=1}^{n}K(\bm{p}_{i},\bm{p}_{j},\bm{y}_{i},\bm{y}_{j})}.$ (37)
The proof is completed by plugging in the kernel $K_{SA}$. ∎
Proof of Proposition III.1
Proof. Using the definition in Definition III.1, we have
$\geq\frac{\|\bm{u}\|^{2}+\|\hat{\bm{u}}\|^{2}+2\|\bm{u}\|\|\hat{\bm{u}}\|\cos(\bm{u}\mid\hat{\bm{u}})}{\|\bm{\eta}\|^{2}+\|\hat{\bm{\eta}}\|^{2}+2\|\bm{\eta}\|\|\hat{\bm{\eta}}\|}\frac{\|\bm{\eta}\|^{2}}{\|\bm{u}\|^{2}}$ (38)
$\geq\frac{(1+\alpha^{2})\|\bm{u}\|^{2}+2\alpha\beta\|\bm{u}\|^{2}}{(1+\gamma)\|\bm{\eta}\|^{2}+2\gamma\|\bm{\eta}\|^{2}}\frac{\|\bm{\eta}\|^{2}}{\|\bm{u}\|^{2}}$ (39)
$=\frac{1+2\alpha\beta+\alpha^{2}}{(1+\gamma)^{2}},$ (40)
$\frac{\sigma(\bm{y}+\hat{\bm{y}})}{\sigma(\bm{y})}\geq\frac{\sqrt{1+2\alpha\beta+\alpha^{2}}}{1+\gamma}>\frac{\sqrt{1+2\gamma+\gamma^{2}}}{1+\gamma}=1$
as desired. ∎
Proof of Theorem III.2
Proof. First, note that the Lipschitz continuity of softmax gives
$\|\mathrm{softmax}(\bm{c}+\bm{\eta})-\mathrm{softmax}(\bm{c})\|\leq L\|\bm{\eta}\|.$ (41)
@@
$\mathbb{E}[\xi]=\mathbb{E}\left[\frac{1}{N}\sum_{i=1}^{N}\eta_{i}^{2}\right]=\frac{1}{N}\sum_{i=1}^{N}\mathbb{E}\left[\eta_{i}^{2}\right]=1,$ (42)
@@
$\Pr(\left|\xi-1\right|\leq\varepsilon)\geq 1-\frac{1}{N\varepsilon^{2}}.$ (43)
@@
$1+\frac{\xi-1}{2}-\frac{(\xi-1)^{2}}{\gamma}\leq\sqrt{\xi}\leq 1+\frac{\xi-1}{2}.$ (44)
@@
Note that since the inequalities applied after the Lipschitzness of softmax are asymptotically tight in $N$, the final bound is also asymptotically tight. This completes the proof. ∎
Proof of Proposition IV.1
Suppose the assumptions described in Remark III.2 hold. Then, the amount of salient information captured by $\bm{y}_{\ell+1}$ is given by
$s_{\ell}:=\alpha^{\ell}\|\bm{u}_{0}\|+\alpha^{i_{\ell}}\|\bm{u}_{0}\|=\alpha^{i_{\ell}}(\alpha^{\ell-i_{\ell}}+1)\|\bm{u}_{0}\|.$ (45)
$\lim_{\ell\to\infty}s_{\ell}=\lim_{\ell\to\infty}\alpha^{i_{\ell}}(\alpha^{\ell-i_{\ell}}+1)\|\bm{u}_{0}\|=0.$ (46)
Therefore, it is necessary for the sequence $\{i_{\ell}\}_{\ell\in\mathbb{N}}$ to contain at least one bounded subsequence to avoid token uniformity. ∎
-A Proof of Theorem IV.1
@@
$\|\bm{Y}_{\ell+1}-\bm{Y}_{\ell+1}^{\prime}\|$
$\leq L_{\ell}\|\bm{Y}_{\ell}-\bm{Y}_{\ell}^{\prime}\|+t\|\bm{Y}_{0}-\bm{Y}_{0}^{\prime}\|+(1-t)\|\bm{Y}_{\ell}-\bm{Y}_{\ell}^{\prime}\|$
@@
$K_{\ell+1}=[L_{\ell}+(1-t)]K_{\ell}+t,\quad K_{1}=L_{0}+1.$ (49)
Assuming $L_{\ell}=L$, define $a=L+(1-t)$, $b=t$, and solve the recurrence $K_{\ell+1}=aK_{\ell}+b$. The closed-form solution for $a\neq 1$ is then given by
$K_{n}=(K_{1}-\tfrac{b}{a-1})a^{n-1}+\tfrac{b}{a-1}.$ (50)
Thus, the dominant term for large $n$ is $K_{\text{GRC}}\asymp[L+(1-t)]^{n}$. Observe that for any $t\in(0,1]$, we have $L+(1-t)<L+1$, implying slower exponential growth in the modified connection as
@@
and a better robustness. ∎

### v2

@@ 5.3 Vision Tasks on ImageNet-1k and ADE20k
Table 4: Image Segmentation on ADE20k
@@
Figure 4: Object Classification on ImageNet-1k under FGSM and PGD attacks with increasing perturbation budgets (severity level $\times$ 1/255). The figures show that models that employ boosting by GRC are consistently more robust than their counterparts without GRC. The vertical axis is in log-scale.
Figure 5: Left to Right: Input, DeiT [touvron2021training], and Deit-Bilateral (ours). Attention heatmaps averaged over attention heads for DeiT [touvron2021training] and DeiT-Bilateral. It can be observed that the heatmaps produced by DeiT-Bilateral captures extended meaningful regions of input samples.
Empirical analysis. Additionally, in Figure 5, we visualize attention heatmaps of the last layer, derived from the normalized attention weights between the class token and the image patches averaged over heads, for both DeiT with $K_{SA}$ and DeiT Bilateral with $K_{BSA}$. We note that the latter attains more coverage of semantically important objects within the image sample. As we have tested on numerous samples, we hypothesize that this behavior, when happens, is partially influenced by the presence of two additional “rough” cross-terms (token-to-position and position-to-token) in the standard self-attention mechanism. As a result, most heads tend to focus on the region with the strongest positive signal, which can dominate these cross-term interactions, while $K_{BSA}$ is free of this distraction. Figure 6, on the other hand, shows the layer-wise average cosine similarities of the Transformer block output tokens, supporting the claim of Proposition 4.1.
11footnotetext: Given the purpose of LRA, required attention span (RAS) serves as a proxy for how difficult a task is for Transformer-based models. The numbers are approximate copies of Figure 2 in [tay2020long].
@@
6 Discussion on Framework Expressivity and Related Work
6.1 Attention Mechanisms and PE
The image filtering perspective provided in this work effectively builds upon and corrects similar notions initially taken by [nguyen2023mitigating] and then [abdullaev2025transformer] that overlook the presence of positional information. Consequently, the framework becomes more expressive and interpretable as we shall discuss next.
Notably, the empirical observations reported in [ke2021rethinking] align closely with our theoretical perspective. The authors of [ke2021rethinking] identify uniform and uninformative patterns in token-to-position proximity logits, suggesting the need for independent transformations of positional vectors to mitigate their influence. Essentially, their empirical observations resembles the Bilateral Attention formulation as the following form of logits are employed:
(35) $\frac{\bm{x}^{\top}_{i}\bm{W}_{Q}^{\top}\bm{W}_{K}\bm{x}_{j}}{\sqrt{2d}}+\frac{\bm{p}^{\top}_{i}\bm{H}_{Q}^{\top}\bm{H}_{K}\bm{p}_{j}}{\sqrt{2d}}+b_{ij},$
where $\bm{H}_{Q}$ and $\bm{H}_{K}$ are independent learnable projections exclusively for PE, resembling the usage of different bandwidth for BF (Section 3.2). Furthermore, Transformers with no positional encodings have also been observed to produce performance on par with absolute, relative and learnable positional encoding methods with somewhat better length generalization properties as studied by [haviv2022transformer]. Our framework provides a theoretical justification for these empirical observations through the lens of NLM filtering (Section 3.2), where we connect no positional encoding to a bilateral attention with infinitesimal weights given to positional weights.
Beyond offering a unifying perspective on self-attention, our work underscores the potential for leveraging insights from classical image processing techniques to inspire novel attention mechanisms. For instance, Elliptical Attention, a recently proposed variant in [nielsen2024elliptical], computes attention scores as
(36) $\alpha_{ij}=\mathrm{softmax}\left(\frac{\bm{q}_{i}^{\top}\bm{M}\bm{k}_{j}}{\sqrt{d}}\right),$
where $\bm{M}$ is a diagonal matrix with $M_{ii}$ representing the average function variability along the feature dimension $i$. This can be interpreted through the lens of anisotropic nonlocal means denoising techniques [Maleki2013Anisotropic] and adaptive filtering methods [takeda2007lark], which construct spatially varying, non-spherical receptive fields to enhance feature preservation. Similarly, Twicing Attention [abdullaev2025transformer] incorporates a residual correction technique from nonparametric regression, reminiscent of the twicing method in image processing, which our generalized framework naturally encompasses when additional computation is permitted as depicted in the following chart:
@@ 6.1 Attention Mechanisms and PE
6.2 What about MLP blocks?
MLPs are usually not straightforward to interpret. The interpretability of MLPs often requires them to be wide-and-sparse [yang2025moeX]. Therefore, we shall discuss the sparse mixture of experts (SMoE) layer employed in most of the current state-of-the-art models [shazeer2017outrageously, fedus2022switch], conjecturing that its computation may bear a resemblance to techniques of sparse coding that are often utilized in k-SVD denoising [romano2015boosting].
In a SMoE layer, $M>1$ MLPs
$f_{j}(\bm{x})=\bm{P}_{j}\sigma(\bm{Q}_{j}\bm{x}),$
known as experts, are used instead of a single dense one to enhance the model capacity. Only a few of them are then activated for each input token conditionally. To be more precise, given an input token $\bm{x}\in\mathbb{R}^{d}$, the output of the SMoE layer is given by
(37) $\bm{y}=\sum_{j=1}^{M}g_{j}(\bm{x},\bm{\theta})f_{j}(\bm{x})=\sum_{j=1}^{M}g_{j}(\bm{x},\bm{\theta})\bm{P}_{j}\sigma(\bm{Q}_{j}\bm{x}),$
@@ 6.2 What about MLP blocks?
(38) $g_{j}(\bm{x},\bm{\theta})=\begin{cases}\mathrm{softmax}(\bm{x}^{\top}\bm{\theta}_{j}),&\text{if }j\in\text{Top-}k(\bm{x}^{\top}\bm{\theta}_{i})\\ 0,&\text{otherwise}.\end{cases}$
Intuitively, the router functions as a means to cluster the input tokens by assigning groups of tokens to different experts by a similarity measure. This aspect of SMoE architecture has indeed been explored by a recent line of research [dikkala2023on, nielsen2025tight]. We also recall that clustering is a key technique behind most image segmentation algorithms as well [coleman1979image, dhanachandra2015image, yasir2020review].
Now to explore the conjecture mentioned in the beginning of this section, we can rewrite the output in a large matrix form, which highlights the sparse and structured nature of the computation, potentially mirroring the sparse coding principles [aharon2006ksvd] used in k-SVD denoising [romano2015boosting]. Following [liu2023sparseFFN], let us define the intermediate output of each expert $j$ as $\bm{z}_{j}=\sigma(\bm{Q}_{j}\bm{x})\in\mathbb{R}^{k^{\prime}}$, where $k^{\prime}$ is the dimension of the intermediate representation. The contribution of expert $j$ to the output $\bm{y}$ is then $g_{j}(\bm{x},\bm{\theta})\bm{P}_{j}\bm{z}_{j}$, and the total output is the sum of these contributions over all $M$ experts.
Now, define $\bm{D}$ as the horizontal concatenation of the output transformation matrices $\bm{P}_{j}$ for all $M$ experts $\bm{D}=[\bm{P}_{1},\bm{P}_{2},\dots,\bm{P}_{M}]\in\mathbb{R}^{d\times(Mk)}$, where each $\bm{P}_{j}\in\mathbb{R}^{d\times k^{\prime}}$, so $\bm{D}$ has $d$ rows and $Mk^{\prime}$ columns. Next, construct a vector $\bm{z}\in\mathbb{R}^{Mk^{\prime}}$ by vertically stacking the scaled intermediate outputs $g_{j}(\bm{x},\bm{\theta})\bm{z}_{j}$ as
$\bm{z}=\begin{bmatrix}g_{1}(\bm{x},\bm{\theta})\bm{z}_{1}&g_{2}(\bm{x},\bm{\theta})\bm{z}_{2}&\dots&g_{M}(\bm{x},\bm{\theta})\bm{z}_{M}\end{bmatrix}^{\top}$.
@@
With these definitions, the SMoE output can be expressed as a single matrix-vector product
$\bm{y}=\bm{D}\bm{z}$
by construction. The matrix $\bm{D}$ can be viewed as a large “dictionary” of transformations, with $Mk$ columns, while the sparsity of $\bm{z}$ ensures that only a small subset of these transformations (corresponding to the $k$ active experts) is used for each input token. We conjecture that this formulation may be related to k-SVD denoising [aharon2006ksvd, romano2015boosting], where a noisy signal is approximated using a sparse linear combination of dictionary atoms to produce a denoised output.
Table 5: Forward Pass Efficiency in terms of FLOPs (GMac)
@@
7 Concluding Remarks
In this work, we developed a unifying image processing framework to better understand the feature-learning mechanism behind architectural components of Transformer-based models. In addition to addressing the frequently asked question of why positional encodings (PE) are added rather than concatenated in an investigative approach, we also explore how properly incorporating positional information enhances the accuracy and long-range capacity of the self-attention mechanism. Furthermore, we highlight the crucial role of residual connections in improving the model’s fidelity to its input and, as a result, robustness against adversarial attacks.
Future directions. Despite the growing interest in drawing inspiration from classical image processing for designing Transformer architectures [nguyen2023mitigating, abdullaev2025transformer], several powerful techniques remain underexplored [takeda2007lark, kheradmand2013general, buhlmann2003boosting, romano2017little], and yet to be systematically transferred to Transformers or alternative models. Similarly, fast approximations of the bilateral filter [Paris2006], which enable efficient edge-preserving smoothing in real-time applications, could have been considered as a foundation for designing computationally efficient self-attention mechanisms or linear recurrent models.
Limitations. A specific limitation of Bilateral Attention proposed in this work is that it introduces an additional training-time computation complexity (note it does not pose any additional inference cost since positional embeddings are input-independent and can be precomputed after training, see Table 5), which is undesirable under specific resource restricted scenarios. Moreover, a room of ambiguity still remains when it comes to understanding MLP blocks (if possible) other than the attempts in the previous section, which is an interesting open question for further exploration.
Appendix A Technical Proofs
Proof of Theorem 3.1
The proof is similar to the derivation of Eqn. 12 but with vector valued intensity values. The signal $u(\bm{p})\in\mathbb{R}^{d}$ at position $\bm{p}$ is estimated using a (nonparametric) point estimation framework. Specifically, this involves solving a weighted least squares problem:
(39) $\hat{u}(\bm{p}_{i})=\arg\min_{u(\bm{p}_{i})}\mathcal{J}_{K}(u(\bm{p}_{i})),$
@@ Appendix A Technical Proofs
(42) $\hat{u}(\bm{p}_{i})=\frac{\sum_{j=1}^{n}\bm{y}_{j}K(\bm{p}_{i},\bm{p}_{j},\bm{y}_{i},\bm{y}_{j})}{\sum_{j=1}^{n}K(\bm{p}_{i},\bm{p}_{j},\bm{y}_{i},\bm{y}_{j})}.$
The proof is completed by plugging in the kernel $K_{SA}$.
Proof of Theorem 3.2
First, note that the Lipschitz continuity of softmax gives
(43) $\|\mathrm{softmax}(\bm{c}+\bm{\eta})-\mathrm{softmax}(\bm{c})\|\leq L\|\bm{\eta}\|.$
@@
(44) $\mathbb{E}[\xi]=\mathbb{E}\left[\frac{1}{N}\sum_{i=1}^{N}\eta_{i}^{2}\right]=\frac{1}{N}\sum_{i=1}^{N}\mathbb{E}\left[\eta_{i}^{2}\right]=1,$
@@
(45) $\Pr(\left|\xi-1\right|\leq\varepsilon)\geq 1-\frac{1}{N\varepsilon^{2}}.$
@@
(46) $1+\frac{\xi-1}{2}-\frac{(\xi-1)^{2}}{\gamma}\leq\sqrt{\xi}\leq 1+\frac{\xi-1}{2}.$
@@
Note that since the inequalities applied after the Lipschitzness of softmax are asymptotically tight in $N$, the final bound is also asymptotically tight. This completes the proof.
A.1 Proof of Theorem 4.1
@@
$\|\bm{Y}_{\ell+1}-\bm{Y}_{\ell+1}^{\prime}\|$ $\leq L_{\ell}\|\bm{Y}_{\ell}-\bm{Y}_{\ell}^{\prime}\|+t\|\bm{Y}_{0}-\bm{Y}_{0}^{\prime}\|+(1-t)\|\bm{Y}_{\ell}-\bm{Y}_{\ell}^{\prime}\|$
@@
(49) $K_{\ell+1}=[L_{\ell}+(1-t)]K_{\ell}+t,\quad K_{1}=L_{0}+1.$
Assuming $L_{\ell}=L$, define a=L+(1t)a=L+(1-t), $b=t$, and solve the recurrence $K_{\ell+1}=aK_{\ell}+b$. The closed-form solution for $a\neq 1$ is then given by
(50) $K_{n}=(K_{1}-\tfrac{b}{a-1})a^{n-1}+\tfrac{b}{a-1}.$
Thus, the dominant term for large $n$ is $K_{\text{GRC}}\asymp[L+(1-t)]^{n}$. Observe that for any $t\in(0,1]$, we have L+(1t)<L+1L+(1-t)<L+1, implying slower exponential growth in the modified connection as
@@
and a better robustness.
Acknowledgments
We would like to acknowledge the assistance of volunteers in putting together this example manuscript and supplement.